OpenStudy (anonymous):

Factor Completely: x2 + 16

OpenStudy (anonymous):

i'm assuming you mean x^2, which makes the roots to this imaginary. what level are you at? gradeschool? then you may have been asked to solve x^2-16. that's a difference of squares. hint: (x-2)(x+2) expands to x^2-2x+2x-4=x^2-4 (note: 4=2^2)

OpenStudy (radar):

Results in complex factors: (x+4i)(x-4i)